OpenStudy (anonymous):

A plane is flying with an airspeed of 160 miles per hour and heading 145°. The wind currents are running at 25 miles per hour at 175° clockwise from due north. Use vectors to find the true course and ground speed of the plane. (Round your answers to the nearest ten for the speed and to the nearest whole number for the angle.) ( )mi/hr with heading( )

OpenStudy (anonymous):

160cos(145)+25cos(-85)= x coordinate of the plane 160sin(145)+25sin(-85)= y coordinate x^2+y^2= c^2 c is the magnitude

OpenStudy (anonymous):

For c I get 77.99 what do I do after?

OpenStudy (anonymous):

that's your answer

OpenStudy (anonymous):

 160 cos145 gave the x coordinate of the plan, added to the x coordinate of the the wind will give you the truth course. Same with y. Pythagorean theorem would give you the magnitude, the truth speed of the plan with the wind going against the plan.